# Towards a Reward-Free Reinforcement Learning Framework for Vehicle Control

### v1

@@ Abstract
Reinforcement learning plays a crucial role in vehicle control by guiding agents to learn optimal control strategies through designing or learning appropriate reward signals. However, in vehicle control applications, rewards typically need to be manually designed while considering multiple implicit factors, which easily introduces human biases. Although imitation learning methods does not rely on explicit reward signals, they necessitate high-quality expert actions, which are often challenging to acquire. To address these issues, we propose a reward-free reinforcement learning framework (RFRLF). This framework directly learns the target states to optimize agent behavior through a target state prediction network (TSPN) and a reward-free state-guided policy network (RFSGPN), avoiding the dependence on manually designed reward signals. Specifically, the policy network is learned via minimizing the differences between the predicted state and the expert state. Experimental results demonstrate the effectiveness of the proposed RFRLF in controlling vehicle driving, showing its advantages in improving learning efficiency and adapting to reward-free environments.
@@ I Introduction
In the field of vehicle control, although traditional control methods such as Model Predictive Control perform well in certain scenarios, they often struggle to achieve ideal control effects when facing dynamically changing environments [1, 2, 3]. In contrast, Reinforcement Learning (RL), as a learning-based decision-making method, can learn the optimal policy through the interaction between the agent and the environment [4, 5, 6]. The advantage of RL lies in its ability to adapt to the uncertainty and dynamic changes of the environment, learning effective behavioral policies in specific environments through a continuous trial-and-error process [7]. Moreover, RL algorithms can automatically adjust and optimize control parameters through the interaction between the agent and the environment [8]. Therefore, researchers are actively exploring various reinforcement learning algorithms and their applications in vehicle control [9, 10].
@@
In recent years, reinforcement learning has achieved many remarkable accomplishments in the field of vehicle driving control [11, 12, 13]. However, traditional reinforcement learning algorithms, such as the MAD-ARL method [14] and the PETS-MPPI method [15], typically require explicit reward signals to guide the vehicle in learning driving policies. Yet, the design of driving rewards for vehicles must consider many implicit factors, making it difficult to define an appropriate reward function for RL [16, 17]. Furthermore, in some real-world scenarios, reward signals may be very sparse or inaccessible [18, 19].
In existing work, Kamienny et al. [20] propose the Import algorithm, which regularizes the training of recurrent neural network strategies by using informative policies. Import effectively reduces the reliance on explicit rewards, but it still needs to use artificially selected task description information to learn the reward function and then guide the learning of the agent. Zintgraf et al. [21] propose the Varibad algorithm, a Bayesian adaptive deep reinforcement learning method based on meta-learning. It uses a variational auto-encoder (VAE) to learn a low-dimensional random latent representation of the task, enabling the agent to effectively explore and exploit in an unknown environment, but it not be able to learn effective strategies for environments where rewards are completely missing. Hejna et al. [22] propose an Inverse Preference Learning (IPL) algorithm aimed at addressing some issues in traditional preference RL methods by eliminating the need for explicit reward functions. IPL enhances learning stability by simultaneously learning an implicit reward that aligns with expert preferences and the optimal policy. Although IPL does not require an explicit reward function, it still necessitates the human design of an implicit reward for optimization, and the specific form of the implicit reward limits its flexible application. In the case of sparse rewards, Liu et al. [23] propose the Dream algorithm, which learns to automatically identify task-relevant information by leveraging goals and restores this information by learning an exploration objective. It performs well in handling sparse reward issues but requires additional policies to ensure effective exploration when rewards are completely missing.
@@
Additionally, when rewards are completely unknown and reward functions are challenging to design, imitation learning emerges as a viable alternative. Imitation learning observes and learns from the behavior of experts, directly extracting behavioral patterns from demonstration data, thus bypassing the reliance on explicit reward signals. Xiao et al. [24] propose an imitation learning method that enables agents to acquire high-quality behavioral policies by imitating expert behaviors and learning from experience distributions rather than reward functions. Generative Adversarial Imitation Learning (GAIL) [25] draws inspiration from Generative Adversarial Network (GAN) [26], training a discriminator to distinguish between expert trajectories and policy-generated trajectories while optimizing the policy to deceive the discriminator. However, imitation learning typically relies on high-quality expert demonstration data, which includes not only state transitions but also the corresponding action information. In some practical applications, obtaining such action information can be challenging because it may involve privacy protection issues or confidentiality of expertise. Generative Adversarial Imitation from Observation (GAIfO) [27], as a representative method of Imitation from Observation (IfO) [28], builds upon GAIL and requires only expert state trajectories as input, without the need for expert actions. However, its training process requires fine-tuning of the balance between the two components, making it prone to training instability or even collapse. We propose a novel framework that incorporates a target state prediction network, allowing the policy network to be optimized gradually without the need to simultaneously handle the dynamic balance in adversarial training, thereby improving learning efficiency and stability.
Inspired by this, we propose a reward-free vehicle control policy learning method, which updates policies using predicted states and expert states, avoiding reliance on manually designed reward functions. Through this approach, the agent focuses on transitioning from the current state to the target state, rather than maximizing manually designed rewards through trial and error, thereby improving learning efficiency and reducing the influence of improper reward function design. However, this reward-free learning also presents new challenges. The agent needs to employ alternative methods to evaluate and optimize its behavior while maintaining the direction of learning in the absence of immediate reward feedback.
Therefore, driven by the insights mentioned above, we propose the RFRLF framework. The RFRLF is mainly divided into two parts: the Target State Prediction Network (TSPN) and the Reward-Free State-Guided Policy Network (RFSGPN). The TSPN aims to predict the state of the environment at next time step based on the agent’s current state and the actions it performs, thereby providing target state predictions for the optimization of subsequent policy networks. The RFSGPN aims to make optimized action decisions without relying on manually designed reward signals. Without the TSPN, the interaction between the agent and the environment cannot perform backpropagation when training the RFSGPN. The TSPN, combined with the state information provided by expert data, provides a loss for policy optimization for the RFSGPN. By training the policy network using a loss function based on state representation, there is no need to depend on manually designed reward signals. Moreover, RFRLF is particularly suitable for situations where expert actions are unknown. By learning target states rather than specific actions, RFRLF enables agents to optimize their behavioral policies even in the absence of expert action guidance, through expert state information. Experimental results demonstrate the effectiveness of the proposed framework in controlling vehicles.
@@
    We propose a novel reward-free reinforcement learning framework (RFRLF). RFRLF can optimize agent behavior by solely observing records (e.g., videos) of expert demonstrations without the need for known rewards or the design of implicit rewards.
@@
    We innovatively combine a TSPN with a RFSGPN, enabling efficient policy optimization without the need to design a reward function, thus mitigating performance issues that may arise from improper reward design.
@@
    This method is particularly suitable for situations where expert actions are unknown, allowing agents to learn effective policies through imitation learning even in the absence of action guidance.
@@ II-A Reinforcement and Imitation Learning in Vehicle Control
In [35], a method called S2RL is proposed. This method uses Control Barrier Functions (CBFs) and Control Lyapunov Functions (CLFs) to guide the training of the vehicle, ensuring that the learned policy maintains optimal behavior while meeting safety and stability requirements. In [36], the authors built the DDQN framework by introducing a dual network architecture, proposing the Dueling Double Deep Q-Network (D3DQN). The D3DQN has shown good performance in both simulation and real-time testing, effectively navigating vehicles around temporary obstacles and pedestrians. However, these methods typically require explicit reward signals to guide the vehicle’s learning to optimize its driving performance. In contrast, our proposed RFRLF method can guide the vehicle to effectively learn driving policies without providing manually designed reward signals. In [37], a Life-long Policy Learning (LLPL) framework is proposed. This framework combines Imitation Learning (IL) with Life-long Learning (LLL), providing vehicles with a path tracking control policy that can continuously learn and adapt to new environments, significantly improving the learning efficiency and adaptability of the policy. In [38], a method based on imitation learning is proposed, which effectively improve the planning and control performance of the vehicle by combining data augmentation, task loss, attention mechanisms, and post-processing planners, achieving robust driving by imitating human driver behavior. However, these methods all require explicit action information in expert data to guide the vehicle’s learning of driving policies and optimize performance. In contrast, our proposed RFRLF method learn effective driving policies without providing explicit action information.
II-B Reinforcement Learning with No Rewards
@@ III Methodology
This chapter provides a detailed introduction to the reward-free reinforcement learning framework (RFRLF) we propose, as shown in Fig. 1. The motivation of this framework stems from addressing the challenge of dependency on reward signals in traditional reinforcement learning. In practical applications, crafting effective reward functions is often difficult and time-consuming, especially in environments where rewards are sparse or ambiguous. To tackle this, we divide the framework into two complementary modules: the Target State Prediction Network (TSPN) and the Reward-Free State-Guided Policy Network (RFSGPN). The TSPN predicts the next state based on the agent’s current state and actions, providing a clear target information for the policy network. The RFSGPN utilizes the target state predictions and expert state information to optimize the agent’s policy, enabling effective learning in environments without explicit rewards. These two modules work in tandem; the TSPN supplies the necessary target information, and the policy network uses this information to refine actions, thereby reducing reliance on manually designed reward signals and enhancing adaptability and learning efficiency in complex environments. Distinguished from neural network-based supervised control methods [42, 43], the proposed method offers the following advantages:
@@ III-A Target State Prediction Network (TSPN)
The design of the TSPN stems from an in-depth understanding of environmental dynamics, aiming to predict the state at the next time step by modeling the interactions between the agent and the environment, thereby providing predictions of target states for subsequent policy optimization. This approach circumvents the reliance on manually designed reward signals in traditional reinforcement learning, enabling the policy network to directly learn how to achieve target states. This network uses random interaction data $(s_{t},a_{t})$ from the agent and the environment to predict the target state $s_{t+1}^{pre}$, and the predicted target state is used alongside the next state $s_{t+1}$ from the interaction data to optimize the prediction network. The optimized prediction network is then used to train the RFSGPN, thereby assisting the agent in learning the policy. This part of the content will be introduced in the following section III-B.
@@
To provide the TSPN with continuous and diverse state-action pair observation data, we design a data collector and a data processor. As shown in Fig. 1, the data collector uses an untrained policy network to interact randomly with the environment to obtain observational data. During this process, a series of state-action-next state triplet data generated are saved to the data buffer. Subsequently, the data processor processes the data in the data buffer to ensure that the extracted observation sequences are continuous and do not contain end flags. This process ensures the quality of the training data for the TSPN, enabling it to learn on a continuous and diverse state-action sequence.
@@
Based on the high-quality training data, the TSPN takes the current state and action information as inputs and outputs the prediction of the state at the next time step. As shown in Fig. 2, the TSPN is composed of multiple modules, including the input embedding layer, downsampling layer, upsampling layer, action injection layer, and the output layer. These modules not only effectively capture the features of the current state but also consider the agent’s decisions and actions through the action injection layer, enabling the network to better capture the correlation between the previous and subsequent states as well as the actions. In the following sections, we will provide a detailed introduction to these modules.
@@
The input embedding layer uses a convolutional neural network to embed the input state frames. We have selected appropriate kernel sizes and strides to ensure that key information in the state frames can be effectively extracted, thereby improving the accuracy of predictions [44]. We design two downsampling layers that reduce the dimension of features obtained after embedding through multiple convolutions and instance normalization operations, reducing computational complexity and memory consumption and allowing the network to consider contextual information over a larger range [45, 46]. Specifically, after the input state $s_{t}$ is embedded and downsampled, the resulting feature representation is:
@@
Figure 2: The structure of the TSPN. In these networks, the input state is first processed by the input layer and then passed through the downsampling layer to extract key state features. Further, the action information is fused into these state features through the action injection layer. Finally, through the upsampling operation and the output layer, the network generates a prediction of the target state.
where $W_{in}$, $b_{in}$, $W_{c1}$, $b_{c1}$, $W_{c2}$ and $b_{c2}$ are the weight matrices and bias vectors of the network, respectively. On this basis, the action injection layer injects action information into the feature representation, enabling the network to make state predictions based on action information. In reinforcement learning tasks, the next state depends not only on the current state but also on the actions taken [47]. Therefore, injecting action information during the downsampling and upsampling processes of the network can help the network better understand the impact of actions on state changes, thereby improving the accuracy of predictions. The action injection layer converts the action vector into a mask through a fully connected layer and performs element-wise multiplication and addition operations with the state features. The specific process is as follows:
@@
where $a_{t}$ is the current action, $W_{a1}$, $W_{a2}$ are the weight matrices of the action injection layer, $m_{1}(t)$, $m_{2}(t)$ are the generated masks, $\sigma$ is the sigmoid function, and $\odot$ represents element-wise multiplication. Subsequently, the upsampling layer upsamples the feature representation through multiple deconvolutions and instance normalization operations, generating a prediction result of the same size as the original input. We designed two upsampling layers to map the features after action injection layer, $h_{2}^{{}^{\prime\prime}}(t)$, back to a higher-dimensional representation $h_{1}^{{}^{\prime}}(t)$ and the final embedded representation $h_{0}^{{}^{\prime}}(t)$ in sequence:
@@
Finally, the output layer uses a convolutional neural network to map the upsampled feature representation to the predicted result of the next state. This design allows the network to directly output pixel-wise prediction results, which is essential for image prediction tasks [48]. The predicted state expression after the output layer is:
@@
It is worth noting that when the input to the TSPN consists of numerical data, convolutional layers are not required for prediction. Therefore, in this case, we replace the original convolutional and deconvolutional layers with linear layers. Additionally, the instance normalization function is replaced by a layer normalization function, which is more appropriate for linear layers.

### v2

@@ Abstract
Reinforcement learning plays a crucial role in vehicle control by guiding agents to learn optimal control strategies through designing or learning appropriate reward signals. However, in vehicle control applications, rewards typically need to be manually designed while considering multiple implicit factors, which easily introduces human biases. Although imitation learning methods does not rely on explicit reward signals, they necessitate high-quality expert actions, which are often challenging to acquire. To address these issues, we propose a reward-free reinforcement learning framework (RFRLF). This framework directly learns the target states to optimize agent behavior through a target state prediction network (TSPN) and a reward-free state-guided policy network (RFSGPN), avoiding the dependence on manually designed reward signals. Specifically, the policy network is learned via minimizing the differences between the predicted state and the expert state, without dependency on expert actions. Experimental results demonstrate the effectiveness of the proposed RFRLF in controlling vehicle driving, showing its advantages in improving learning efficiency and adapting to reward-free environments.
@@ I Introduction
In the field of vehicle control, although traditional control methods such as Model Predictive Control perform well in certain scenarios, they often struggle to achieve ideal control effects when facing dynamically changing environments [1, 2, 3]. In contrast, Reinforcement Learning (RL), as a learning-based decision-making method, can learn the policy through the interaction between the agent and the environment [4, 5, 6]. The advantage of RL lies in its ability to adapt to the uncertainty and dynamic changes of the environment, learning effective behavioral policies in specific environments through a trial-and-error process [7]. Moreover, RL algorithms can automatically adjust and optimize control parameters through the interaction between the agent and the environment [8]. Therefore, researchers are actively exploring various reinforcement learning algorithms and their applications in vehicle control [9, 10].
@@
In recent years, reinforcement learning has achieved many remarkable accomplishments in the field of vehicle driving control [11, 12, 13]. However, traditional reinforcement learning algorithms, such as the MAD-ARL method [14] and the PETS-MPPI method [15], typically require explicit reward signals to guide the vehicle in learning driving policies. Yet, the design of driving rewards for vehicles must consider many implicit factors, making it difficult to define an appropriate reward function for RL [16, 17]. Furthermore, in some real-world scenarios, reward signals are very sparse or inaccessible [18, 19].
Multiple approaches have been proposed to deal with sparse reward signals. Kamienny et al. [20] propose the Import algorithm which utilizes informative policies to regularize recurrent neural network training, reducing dependency on explicit rewards, but still requires manually selected task descriptions to learn reward functions. Zintgraf et al. [21] develop the Varibad algorithm, a Bayesian adaptive approach based on meta-learning, which employs a Variational Auto-Encoder (VAE) to learn low-dimensional latent representations of tasks, yet proves ineffective in environments with completely absent rewards. Hejna et al. [22] introduce the Inverse Preference Learning (IPL) algorithm that enhances stability by simultaneously learning implicit rewards and optimal policies, though it still necessitates manually designed implicit rewards, limiting its application flexibility. Liu et al. [23] presente the Dream algorithm that identifies task-relevant information by learning exploitation objectives, demonstrating excellent performance in sparse reward scenarios, but requiring additional strategies to ensure effective exploration in completely reward-free environments.
@@
Additionally, when rewards are completely unknown and reward functions are challenging to design, imitation learning emerges as a viable alternative. Imitation learning observes and learns from the behavior of experts, directly extracting behavioral patterns from demonstration data, thus bypassing the reliance on explicit reward signals. Xiao et al. [24] propose an imitation learning method that enables agents to acquire high-quality behavioral policies by imitating expert behaviors and learning from experience distributions rather than reward functions. Generative Adversarial Imitation Learning (GAIL) [25] draws inspiration from Generative Adversarial Network (GAN) [26], training a discriminator to distinguish between expert trajectories and policy-generated trajectories while optimizing the policy to deceive the discriminator. However, imitation learning typically relies on high-quality expert demonstration data, which includes not only state transitions but also the corresponding action information. In some practical applications, obtaining such action information can be challenging because it may involve privacy protection issues or confidentiality of expertise. Generative Adversarial Imitation from Observation (GAIfO) [27], as a representative method of Imitation from Observation (IfO) [28], builds upon GAIL and requires only expert state trajectories as input, without the need for expert actions. However, its training process requires fine-tuning the balance between the generator and the discriminator, which can easily lead to training instability or even collapse.
To achieve vehicle control without designing reward functions and without expert actions, we propose a Reward-Free Reinforcement Learning Framework (RFRLF), which updates policies using predicted states and expert states while eliminating dependency on manually designed reward functions. The RFRLF comprises two core components: a Target State Prediction Network (TSPN) and a Reward-Free State-Guided Policy Network (RFSPN). The TSPN predicts the state at the next timestep based on the agent’s current state and executed action, and the predicted states is then used to optimize policy network (i.e., RFSPN) without relying on manually designed reward signals. Without the TSPN, the discrepancy between the agent states and expert states cannot directly guide policy learning through backpropagation during the training phase of the RFSPN due to the unknown environment model. Furthermore, the TSPN enables gradual policy network optimization and eliminates the need to simultaneously handle adversarial training dynamics, thereby enhancing learning stability. By training the policy network through loss functions based on state representations rather than manually crafted rewards, we mitigate the human biases introduced by reward function design. Additionally, the RFRLF demonstrates particular suitability for scenarios with unknown expert actions, which are often necessary in conventional imitation learning and used as a substitute for rewards as supervision signals. Experimental results demonstrate the effectiveness of the proposed framework in controlling vehicles.
@@
    We propose a novel Reward-Free Reinforcement Learning Framework (RFRLF). RFRLF learns agent behavior by observing records of expert demonstrations (such as videos).
@@
    We design target state prediction networks (TSPN) and reward-free state-guided policy networks (RFSGPN) for the proposed framework. Through their tandem update, policies can be efficiently optimized without reward function design, mitigating performance issues caused by improper reward engineering.
@@
    This method is particularly suitable for situations where expert actions are unknown, allowing agents to learn effective policies through imitation learning even in the absence of expert actions.
@@ II-A Reinforcement and Imitation Learning in Vehicle Control
In [35], a method called S2RL is proposed. This method uses Control Barrier Functions (CBFs) and Control Lyapunov Functions (CLFs) to guide the training of the vehicle, ensuring that the learned policy maintains optimal behavior while meeting safety and stability requirements. In [36], the authors built the DDQN framework by introducing a dual network architecture, proposing the Dueling Double Deep Q-Network (D3DQN). The D3DQN has shown good performance in both simulation and real-world experiments, effectively navigating vehicles around temporary obstacles and pedestrians. However, these methods typically require explicit reward signals to guide the vehicle’s learning to optimize its driving performance. In contrast, our proposed RFRLF method can guide the vehicle to effectively learn driving policies without providing manually designed reward signals. In [37], a Life-long Policy Learning (LLPL) framework is proposed. This framework combines Imitation Learning (IL) with Life-long Learning (LLL), providing vehicles with a path tracking control policy that can continuously learn and adapt to new environments, significantly improving the learning efficiency and adaptability of the policy. In [38], a method based on imitation learning is proposed, which effectively improve the planning and control performance of the vehicle by combining data augmentation, task loss, attention mechanisms, and post-processing planners, achieving robust driving by imitating human driver behavior. However, these methods all require action information in expert data to guide the vehicle’s learning of driving policies and optimize performance. In contrast, our proposed RFRLF method learn effective driving policies without providing action information of experts.
II-B Reinforcement Learning without Rewards
@@ III Methodology
In practical applications, designing effective reward functions is often challenging and time-consuming, especially in environments with sparse or ambiguous rewards. To address the challenge of reward signal dependency in traditional reinforcement learning, this chapter details our proposed Reward-Free Reinforcement Learning Framework (RFRLF), as illustrated in Fig. 1. We divide the framework into two complementary modules: the Target State Prediction Network (TSPN) and the Reward-Free State-Guided Policy Network (RFSGPN). The TSPN predicts the next state based on the agent’s current state and actions, providing a clear target information for the policy network. The RFSGPN utilizes the target state predictions and expert state information to optimize the agent’s policy, enabling effective learning in environments without explicit rewards. These two modules work in tandem; the TSPN supplies the necessary target information, and the policy network uses this information to refine actions, thereby reducing reliance on manually designed reward signals and enhancing adaptability and learning efficiency in complex environments. Distinguished from neural network-based supervised control methods [42, 43], the proposed method offers the following advantages:
@@ III-A Target State Prediction Network (TSPN)
The design of the TSPN stems from an in-depth understanding of environmental dynamics, aiming to predict the state at the next time step by modeling the interactions between the agent and the environment, thereby providing predictions of target states for subsequent policy optimization. This approach circumvents the reliance on manually designed reward signals in traditional reinforcement learning, enabling the policy network to directly learn how to achieve target states. This network uses random interaction data $(s_{t},a_{t})$ from the agent and the environment to predict the target state $s_{t+1}^{pre}$, and the predicted target state is used alongside the next state $s_{t+1}$ from the environmental interaction to optimize the prediction network. The optimized prediction network is then used to train the RFSGPN, which serves as the policy network. This part of the content will be introduced in the following section III-B.
@@
To provide the TSPN with continuous and diverse state-action pair observation data, we design a data collector and a data processor. As shown in Fig. 1, the data collector uses an untrained policy network to interact randomly with the environment to obtain observational data. During this process, a series of (state, action, next state) triplet data generated are saved to the data buffer. Subsequently, the data processor processes the data in the data buffer to ensure that the extracted observation sequences are continuous and do not contain end flags. This process ensures the quality of the training data for the TSPN, enabling it to learn on diverse state-action sequences.
@@
Based on the high-quality training data, the TSPN takes the current state and action information as inputs and outputs the prediction of the state at the next time step. As shown in Fig. 2, the TSPN is composed of multiple modules, including the input embedding layer, feature extraction layer, feature reconstruction layer, action injection layer (including semantic mapping, mask generation and feature fusion), and the spatial decoding layer. These modules not only effectively capture the features of the current state but also consider the agent’s decisions and actions through the action injection layer, enabling the network to better capture the correlation between the previous and subsequent states as well as the actions. In the following sections, we will provide a detailed introduction to these modules.
Figure 2: The structure of the target state prediction network (TSPN). In these networks, the input state is first processed by the input layer and then passed through the feature extraction layer to extract key state features. Further, the action information is fused into these state features through the action injection layer. Finally, through the feature reconstruction and the spatial decoding layer, the network generates a prediction of the target state.
@@
The input embedding layer uses a convolutional neural network to embed the input state frames. We select appropriate kernel sizes and strides to ensure that key information in the state frames can be effectively extracted, thereby improving the accuracy of predictions. We design two feature extraction layers that reduce the dimension of features obtained after embedding through multiple convolutions and instance normalization operations, reducing computational complexity and memory consumption and allowing the network to consider contextual information over a larger range. Specifically, the input state is processed as follows:
@@
where $W_{in}$, $b_{in}$, $W_{c1}$, $b_{c1}$, $W_{c2}$ and $b_{c2}$ are the weight matrices and bias vectors of the network, respectively.
On this basis, the action injection layer injects action information into the feature representation, enabling the network to make state predictions based on action information. In reinforcement learning tasks, the next state depends not only on the current state but also on the actions taken. Therefore, injecting action information during the feature extraction and feature reconstruction of the network can help the network better understand the impact of actions on state changes, thereby improving the accuracy of predictions. The action injection layer employs dual-branch fully connected layers to generate action masks: the Mask Generation branch maps actions to a channel attention mask $m_{1}(t)$ , whose elements are either 0 or 1, and the Semantic Mapping branch linearly transforms actions into a compensation mask $m_{2}(t)$. During feature fusion, the layer first performs element-wise multiplication between the state feature $h_{2}(t)$ and $m_{1}(t)$, then applies residual compensation by adding $m_{2}(t)$, ultimately producing enhanced features with integrated action information. This process is formulated as follows:
@@
where $a_{t}$ is the current action, $W_{a1}$, $W_{a2}$ are the weight matrices of the action injection layer, $\sigma$ is the sigmoid function, and $\odot$ represents element-wise multiplication. Subsequently, the feature reconstruction layer reconstructs the feature representation through multiple deconvolution and instance normalization operations. We designe two feature reconstruction layers to map the features after action injection layer, $h_{2}^{{}^{\prime\prime}}(t)$, back to a higher-dimensional representation $h_{1}^{{}^{\prime}}(t)$ and the final embedded representation $h_{0}^{{}^{\prime}}(t)$ in sequence:
@@
Finally, the spatial decoding layer uses a convolutional neural network to spatially decode the reconstructed features and obtain the prediction of the next state. This design allows the network to directly output pixel-wise prediction results, which is formulated by
@@
It is worth noting that when the input to the TSPN consists of numerical data, convolutional layers are not required for prediction. Therefore, in this case, we replace the original convolutional and deconvolutional layers with fully-connected layers. Additionally, the instance normalization function is replaced by a layer normalization function, which is more appropriate for fully-connecteds layers.